# Ambient Diffusion Posterior Sampling: Solving Inverse Problems with Diffusion Models trained on Corrupted Data

### v1

@@ Abstract
We provide a framework for solving inverse problems with diffusion models learned from linearly corrupted data. Our method, Ambient Diffusion Posterior Sampling (A-DPS), leverages a generative model pre-trained on one type of corruption (e.g. image inpainting) to perform posterior sampling conditioned on measurements from a potentially different forward process (e.g. image blurring). We test the efficacy of our approach on standard natural image datasets (CelebA, FFHQ, and AFHQ) and we show that A-DPS can sometimes outperform models trained on clean data for several image restoration tasks in both speed and performance. We further extend the Ambient Diffusion framework to train MRI models with access only to Fourier subsampled multi-coil MRI measurements at various acceleration factors (R$=2,4,6,8$). We again observe that models trained on highly subsampled data are better priors for solving inverse problems in the high acceleration regime than models trained on fully sampled data. We open-source our code and the trained Ambient Diffusion MRI models: github.com/utcsilab/ambient-diffusion-mri.
@@ 1 Introduction
For some applications, it is expensive or impossible to acquire fully observed or uncorrupted data [15, 20, 47] but possible to acquire partially observed samples. Further, in some cases, it may be desirable to train generative models with noisy or corrupted data since that reduces memorization of the training set [17, 8, 37]. Prior works have shown how to train Generative Adversarial Networks (GANs) [7, 14], flow models [29] and restoration models  [32, 31, 47, 35] with corrupted training data. More recently, there has been a shift towards training diffusion generative models given corrupted data  [17, 1, 28, 16, 30]. What remains unexplored is how to use generative models trained on a certain type of corruption (e.g. inpainted data) to solve inverse problems that arise from a different forward process (e.g. blurring). Thus far, it is not known whether models trained on corrupted data are useful priors for solving inverse problems.
Figure 1: Illustration of Ambient Diffusion Posterior Sampling (Ambient DPS). During training, we only have access to linearly corrupted data from a forward operator $A_{\mathrm{train}}$. We use this data and the Ambient Diffusion framework to learn a generative model, $G_{\mathrm{ambient}}$, for the uncorrupted distribution, $p({\bm{x}}_{0})$. At inference time, we use the learned generator to sample from the posterior distribution $p({\bm{x}}_{0}|{\bm{y}}_{\mathrm{A_{\mathrm{inf}}}})$, for measurements ${\bm{y}}_{\mathrm{inf}}$ coming from a different forward operator, $A_{\mathrm{inf}}$.
We propose the first framework to solve inverse problems with diffusion models learned from linearly corrupted data, for example as in Ambient Diffusion [17]. Ambient Diffusion models estimate the ambient score, i.e. how to best reconstruct given an input with a corrupted linear forward operator corrupted noisy input. We show how to use these models for solving linear inverse problems outside of their training distribution. Our experiments on standard datasets of natural images and multi-coil MRI show something surprising: Ambient Models can outperform (in the high corruption regime) models trained on clean data. Further, they do so while being substantially faster. Our algorithm extends Diffusion Posterior Sampling [12] to Ambient Models and fully unlocks the potential of generative models trained on corrupted data for image restoration.
Our contributions:
  1. 1.
@@
  2. 2.
    We use pre-trained Ambient Diffusion models to solve inverse problems (compressed sensing, super-resolution) on several natural image datasets (CelebA, FFHQ, AFHQ) and show that they can even outperform models trained on clean data in the high corruption regime.
  3. 3.
    We extend the Ambient Diffusion framework to train models directly on multi-coil Fourier subsampled MRI.
  4. 4.
    We train Ambient Diffusion MRI models on FastMRI at various retrospective acceleration factors (R=$2,4,6,8$); we again observe that models trained on subsampled data are better priors for solving inverse problems in the high acceleration regime.
  5. 5.
    We open-source our code and models to accelerate research in this area: github.com/utcsilab/ambient-diffusion-mri.
@@ 2.1 Background and Notation
Diffusion models are typically trained (up to network reparametrizations) to reconstruct a clean image ${\bm{x}}_{0}\sim p_{0}({\bm{x}}_{0})$ from a noisy observation ${\bm{x}}_{t}={\bm{x}}_{0}+\sigma_{t}\bm{\eta},\quad\eta\sim\mathcal{N}(\mathbf{0},I)$. Despite the simplicity of the training objective, diffusion models can approximately sample from $p({\bm{x}})$ by running a discretized version of the Stochastic Differential Equation [39, 3]:
@@
where ${\bm{w}}$ is the standard Wiener process and $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ is estimated by the trained neural network. Given a measurement ${\bm{y}}_{\mathrm{inf}}=A_{\mathrm{inf}}{\bm{x}}_{0}$, one can sample from the posterior distribution $p({\bm{x}}_{0}|{\bm{y}}_{\mathrm{inf}})$ by running the process:
@@
For most forward operators it is intractable to write the likelihood in closed form. Hence, several approximations have been proposed to use diffusion models for inverse problems [12, 27, 25, 38, 13, 19, 21]. One of the simplest and most effective approximations is Diffusion Posterior Sampling (DPS) [12]. DPS estimates ${\bm{x}}_{0}$ using ${\bm{x}}_{t}$ and uses the conditional likelihood $p({\bm{y}}_{\mathrm{inf}}|\hat{\bm{x}}_{0})$ instead of the intractable term, i.e. DPS approximates $p({\bm{y}}_{\mathrm{inf}}|{\bm{x}}_{t})$ with $p({\bm{y}}_{\mathrm{inf}}|{\bm{x}}_{0}=\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}])$. The update rule becomes:
@@
where $\gamma_{t}$ is a tunable guidance parameter.
@@
[17, ] considered the setting of having access to linearly corrupted data $\{{\bm{y}}_{0}=A_{\mathrm{train}}x_{0},A_{\mathrm{train}}\}$, where the distribution of $A_{\mathrm{train}}$, denoted as $p(A_{\mathrm{train}})$, is assumed to be known. The ultimate goal is to learn the best restoration model given a linearly corrupted noisy input ${\bm{y}}_{t,\mathrm{train}}=A_{\mathrm{train}}({\bm{x}}_{0}+\sigma_{t}\bm{\eta})$, at all noise levels $t$. [17, ] form further corrupted iterates $\tilde{\bm{y}}_{t,\mathrm{train}}=\tilde{A}_{\mathrm{train}}({\bm{x}}_{0}+\sigma_{t}\bm{\eta})$ and train with the following objective:
@@
MRI is a prototypical use case for a framework that can learn generative models from linearly corrupted data, as in many cases it is not feasible to collect a large training set of fully sampled data [43, 18, 42]. In settings such as 3D+time dynamic contrast-enhanced MRI [49] it is impossible to collect fully sampled data due to the time-varying dynamics of the contrast agent [50].
@@
where ${\bm{x}}$ is the (complex-valued) image of interest, $S_{i}$ represents the coil-sensitivity profile of the $i^{th}$ coil, $\mathcal{F}$ is the (2-D or 3-D) Fourier transform, $P$ represents the Fourier subsampling operator and ${\bm{w}}_{i}$ is complex-valued Gaussian i.i.d noise. For simplicity in this work, we assume that we work in the noiseless case and we point the reader to the works of [1, 28, ] for approaches that handle the noisy case.
@@
Due to the time required to collect $k$-space measurements, it is often not possible to acquire fully sampled $k$-space for an entire dataset. This means that training cannot rely on a fully sampled image to guide reconstruction quality as is done in the fully supervised setting [2, 22]. To overcome this limitation, some end-to-end techniques use a loss on the measurement domain by partitioning the training measurements for each sample into (1) measurements for reconstruction, and (2) measurements for applying the loss function [47]. Other approaches leverage structure in the MRI acquisition to learn from limited-resolution data [46]. More recently, works have begun leveraging signal set properties such as group invariance to assist in learning from subsampled measurements for a variety of inverse problems [40, 36, 9, 41].
All these approaches learn a restoration model without access to fully sampled reference data. However, as they are inherently end-to-end methods, their performance on out-of-distribution tasks (e.g., due to different acquisition trajectories) is known to degrade [25, 48].
@@ 2.2 Ambient Diffusion for MRI
As explained above, the MRI acquisition process results in linearly corrupted measurements and thus it should be possible to use the Ambient Diffusion framework to learn diffusion models from the corrupted observations. Yet, we identified three important changes that differ from the setting studied by [17, ]: i) the inpainting happens in the Fourier Domain, ii) the inpainting has structure, i.e. whole vertical lines in the spatial Fourier are either observed or not observed (instead of masking random pixels) and, iii) the image is measured from an array of spatially varying receive coils. In what follows, we show how to account for these factors and apply the Ambient Diffusion framework to MRI data.
@@
Note that this is always possible to do since we can just subsample the available data ${\bm{z}}_{{\bm{x}},i}$. We also underline that to further corrupt the given measurements we keep the inpainting structure that there is in the measurements, i.e. we delete (at random) 2-D lines in the Fourier space (instead of independent pixel masking as in [17, ]).
In that regards the multiplicity of the coils, we combine their measurements to form a crude estimator of ${\bm{x}}$ by taking the adjoint of $\tilde{A}_{\mathrm{train}}$:
$\tilde{\bm{y}}_{\mathrm{train}}=\sum_{i}S_{i}^{H}\mathcal{F}^{-1}(\tilde{\bm{z}}_{{\bm{x}},i}).$ (2.7)
Notice that for a discrete signal ${\bm{x}}\in\mathbb{C}^{n}$, all these operations are linear and hence $\tilde{\bm{y}}$ can be written as:
@@
Let ${\bm{x}}_{t},{\bm{y}}_{t,\mathrm{train}},\tilde{\bm{y}}_{t,\mathrm{train}}$ represent the noisy versions of ${\bm{x}},{\bm{y}}_{\mathrm{train}},\tilde{\bm{y}}_{\mathrm{train}}$ respectively, i.e.:
$\begin{cases}{\bm{x}}_{t}={\bm{x}}+\sigma_{t}\bm{\eta}\\ {\bm{y}}_{t,\mathrm{train}}=A_{\mathrm{train}}\left({\bm{y}}_{t,\mathrm{train}}+\sigma_{t}\bm{\eta}\right)\\ \tilde{\bm{y}}_{t,\mathrm{train}}=\tilde{A}_{\mathrm{train}}\left(\tilde{\bm{y}}_{t,\mathrm{train}}+\sigma_{t}\bm{\eta}\right)\end{cases}.$ (2.10)
@@
The formal proof of this Theorem is given in the Appendix. We provide a sketch of the proof here. The techniques are based on Theorem 4.2 of Ambient Diffusion [17]. In Theorem 4.2, the condition that needs to be satisfied is that $\mathbb{E}[A_{\mathrm{train}}^{T}A_{\mathrm{train}}|\tilde{A}_{\mathrm{train}}]$ is full-rank. Our objective is slightly different because the network doesn’t see the whole forward operator $\tilde{A}_{\mathrm{train}}$, but only part of it, i.e. the matrix $\tilde{P}$. By following the steps of the Ambient Diffusion proof, we arrive at the necessary condition for our case, which is to prove that $\mathbb{E}[A_{\mathrm{train}}|\tilde{P}]$ is full-rank.
@@
Note on alternative Ambient Diffusion designs.
Our approach to extending Ambient Diffusion to the Fourier subsampled MRI case is by no means unique. For example, we could have used other linear operators to aggregate the measurements from the different coils (e.g. we could have used the pseudoinverse), we could further condition on the sensitivity maps, or we could have directly fed all the coil measurements as input to the network without aggregating them. Such approaches could potentially lead to improved performance since they fully leverage the coil information. In this work, we opted for a simple and fast method (aggregation through the adjoint) and we leave such investigations for future work.

### v2

@@ Abstract
We provide a framework for solving inverse problems with diffusion models learned from linearly corrupted data. Firstly, we extend the Ambient Diffusion framework to enable training directly from measurements corrupted in the Fourier domain. Subsequently, we train diffusion models for MRI with access only to Fourier subsampled multi-coil measurements at acceleration factors R$=2,4,6,8$. Secondly, we propose Ambient Diffusion Posterior Sampling (A-DPS), a reconstruction algorithm that leverages generative models pre-trained on one type of corruption (e.g. image inpainting) to perform posterior sampling on measurements from a different forward process (e.g. image blurring). For MRI reconstruction in high acceleration regimes, we observe that A-DPS models trained on subsampled data are better suited to solving inverse problems than models trained on fully sampled data. We also test the efficacy of A-DPS on natural image datasets (CelebA, FFHQ, and AFHQ) and show that A-DPS can sometimes outperform models trained on clean data for several image restoration tasks in both speed and performance. [^0] Equal contribution.
Asad Aali Giannis Daras Brett Levac
Stanford University MIT CSAIL UT Austin
UT Austin IFML IFML
Sidharth Kumar Alexandros G. Dimakis Jonathan I. Tamir
UT Austin UC Berkeley UT Austin
IFML Bespoke Labs IFML
Figure 1: Ambient Diffusion Posterior Sampling (Ambient DPS). During training, we only have access to linearly corrupted data from a forward operator $A_{\mathrm{train}}$. We use the Ambient Diffusion framework to learn a generative model, $G_{\mathrm{ambient}}$, for the uncorrupted distribution, $p({\bm{x}}_{0})$. At inference time, we sample from the posterior distribution $p({\bm{x}}_{0}|{\bm{y}}_{\mathrm{A_{\mathrm{inf}}}})$, for measurements ${\bm{y}}_{\mathrm{inf}}$ coming from a different forward operator, $A_{\mathrm{inf}}$.
Figure 2: Prior samples from diffusion models trained on MRI scans. Columns $1-4$: Diffusion models trained on subsampled MRI scans at acceleration factors $R=2,4,6,8$, using the Ambient Diffusion framework extended for Fourier subsampled training. Columns $5-8$: EDM models trained with L1-wavelet reconstructions of subsampled scans at $R=2,4,6,8$. Column $9$: NCSNV2 trained with fully sampled scans. Column $10$: EDM trained with fully sampled scans. We observe that Ambient Diffusion models consistently produce high-quality and realistic MRI scans even in high acceleration regimes.
@@ 1 Introduction
For some applications, it is expensive or impossible to acquire fully observed data (Collaboration et al., 2019; Gao et al., 2023; Yaman et al., 2020) but possible to acquire partially observed samples. Furthermore, it may be desirable to train generative models with corrupted data since that reduces memorization of the training set (Daras et al., 2023b; Carlini et al., 2023; Somepalli et al., 2022). Prior works have shown how to train Generative Adversarial Networks (GANs) (Bora et al., 2018; Cole et al., 2021), flow models (Kelkar et al., 2023) and restoration models (Lehtinen et al., 2018; Krull et al., 2019; Yaman et al., 2020; Millard & Chiew, 2023) with corrupted data. More recently, there has been a shift towards training diffusion generative models given corrupted data (Daras et al., 2024; 2023a; 2023b; Aali et al., 2023; Kawar et al., 2023; Cui et al., 2022; Kim & Ye, 2021). What remains unexplored is how generative models trained on a certain type of corruption (e.g. inpainted data) can be used to solve inverse problems with a different forward process (e.g. blurring).
We propose the first framework to solve inverse problems with Ambient Diffusion models (Daras et al., 2023b). These models are trained using only access to linear measurements and they estimate the ambient score, i.e. how to best reconstruct given an input with a corrupted linear forward operator corrupted noisy input. We show how to use these models for solving linear inverse problems outside their training distribution. Our experiments on datasets of natural images and multi-coil MRI show something surprising: Ambient Models can outperform (in the high corruption regime) models trained on clean data while being substantially faster. Our algorithm extends Diffusion Posterior Sampling (Chung et al., 2023) to Ambient Diffusion models. Our contributions: [^2] We open-source our code and Ambient Diffusion models: github.com/utcsilab/ambient-diffusion-mri.
00footnotetext:
Figure 3: Posterior sampling reconstructions for MRI scans using models trained on Fourier subsampled data at various acceleration factors (Ambient DPS, columns $2-5$) and a model trained on clean data (FS-DPS (Chung et al., 2023), column $6$). Rows $1$ and $3$ show reconstructions at $R=4$ and $R=8$, respectively, while rows $2$ and $4$ display the difference to the ground truth on a 10×10$\times$10 × scale. At high inference acceleration (R=$8$), Ambient DPS, outperforms FS-DPS, despite that the underlying models were trained solely on corrupted data.
@@
  • We extend the Ambient Diffusion framework to train models using Fourier subsampled measurements. Specifically, we train on subsampled multi-coil MRI scans at various retrospective acceleration factors (R=$2,4,6,8$); we observe that models trained on subsampled data are better priors for solving inverse problems in the high acceleration regime.
  • We use pre-trained Ambient Diffusion models to solve inverse problems (compressed sensing, super-resolution) on natural image datasets (CelebA, FFHQ, AFHQ) and show that they can even outperform models trained on clean data in the high corruption regime.
@@ 2.1 Background and Notation
Diffusion models are typically trained to reconstruct a clean image ${\bm{x}}_{0}\sim p_{0}({\bm{x}}_{0})$ from a noisy observation ${\bm{x}}_{t}={\bm{x}}_{0}+\sigma_{t}\bm{\eta},\quad\eta\sim\mathcal{N}(\mathbf{0},I)$. Despite the simplicity of the training objective, diffusion models can approximately sample from $p({\bm{x}})$ by running a discretized version of the Stochastic Differential Equation (Song et al., 2020; Anderson, 1982):
@@
where ${\bm{w}}$ is the standard Wiener process and $\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]$ is estimated by the network. Given measurement ${\bm{y}}_{\mathrm{inf}}=A_{\mathrm{inf}}{\bm{x}}_{0}$, one can sample from the posterior $p({\bm{x}}_{0}|{\bm{y}}_{\mathrm{inf}})$ by running the process:
@@
For most forward operators it is intractable to write the likelihood in closed form. Hence, several approximations have been proposed to use diffusion models for inverse problems (Chung et al., 2023; Kawar et al., ; Jalal et al., 2021; Song et al., 2021; Chung et al., 2022; Feng et al., 2023; Graikos et al., 2022). One of the simplest and most effective approximations is Diffusion Posterior Sampling (DPS) (Chung et al., 2023). DPS estimates ${\bm{x}}_{0}$ using ${\bm{x}}_{t}$ and uses the conditional likelihood $p({\bm{y}}_{\mathrm{inf}}|\hat{\bm{x}}_{0})$ instead of the intractable term, i.e. DPS approximates $p({\bm{y}}_{\mathrm{inf}}|{\bm{x}}_{t})$ with $p({\bm{y}}_{\mathrm{inf}}|{\bm{x}}_{0}=\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}])$, where $\gamma_{t}$ is a tunable guidance parameter. The update rule becomes:
@@
Figure 4: Posterior sampling reconstructions for FFHQ, showing: (1) Corrupted Input: Examples from FFHQ dataset corrupted using Resolution Downscaling (4×4$\times$4 ×) and Gaussian Noise ($\sigma=0.05$), (2) FS-DPS: Reconstruction using Diffusion Posterior Sampling (DPS), with a diffusion model trained with clean fully-sampled data ($p=0.0$), (3) A-DPS: Reconstruction using Ambient-DPS, with a diffusion model trained on randomly inpainted data with erasure probability ($p=0.6$), (4) Ground Truth: Original uncorrupted examples from FFHQ dataset. We observe that Ambient DPS, provides better reconstructions (qualitatively) even though the underlying models were trained on corrupted data ($p=0.6$).
@@
The authors of Daras et al. (2023b) consider the setting of having access to linearly corrupted data $\{{\bm{y}}_{0}=A_{\mathrm{train}}x_{0},A_{\mathrm{train}}\}$, where the distribution of $A_{\mathrm{train}}$, denoted as $p(A_{\mathrm{train}})$, is assumed to be known. The ultimate goal is to learn the best restoration model given a linearly corrupted noisy input ${\bm{y}}_{t,\mathrm{train}}=A_{\mathrm{train}}({\bm{x}}_{0}+\sigma_{t}\bm{\eta})$, at all noise levels $t$. The authors of Daras et al. (2023b) form further corrupted iterates $\tilde{\bm{y}}_{t,\mathrm{train}}=\tilde{A}_{\mathrm{train}}({\bm{x}}_{0}+\sigma_{t}\bm{\eta})$ using a matrix $\tilde{A}_{\mathrm{train}}$ (that is a perturbation of the given matrix $A_{\mathrm{train}}$) and train with the following objective:
@@
MRI is a prototypical use case for a framework that can learn generative models from linearly corrupted data, as in many cases it is not feasible to collect a large training set of fully sampled data (Tibrewala et al., 2023; Desai et al., 2022; Tariq et al., ). In settings such as 3D+time dynamic contrast-enhanced MRI (Zhang et al., 2015) it is impossible to collect fully sampled data due to the time-varying dynamics of the contrast agent (Zhang et al., ).
@@
where ${\bm{x}}$ is the (complex-valued) image of interest, $S_{i}$ represents the coil-sensitivity profile of the $i^{th}$ coil, $\mathcal{F}$ is the Fourier transform, $P$ represents the Fourier subsampling operator and ${\bm{w}}_{i}$ is complex-valued Gaussian i.i.d noise. We assume the noiseless case and we point the reader to Daras et al. (2023a); Aali et al. (2023); Kawar et al. (2023) for approaches that handle the noisy case.
@@
Due to the time required to collect $k$-space measurements, it is often not possible to acquire fully sampled $k$-space. Hence, training cannot rely on a fully sampled image to guide reconstruction quality as is done in the fully supervised setting (Aggarwal et al., 2019; Hammernik et al., 2018). If two independent measurements are available for the same underlying signal, restoration models can be trained without having access to uncorrupted data (Lehtinen et al., 2018; Gan et al., 2023). If only one measurement is available (as in our setting), some end-to-end techniques use a loss on the measurement domain by partitioning the training measurements for each sample into (1) measurements for reconstruction, and (2) measurements for applying the loss function (Yaman et al., 2020). Other approaches leverage structure in the MRI acquisition to learn from limited-resolution data (Wang et al., 2023). More recently, works have begun leveraging signal set properties such as group invariance to assist in learning from subsampled measurements for a variety of inverse problems (Tachella et al., 2022b; Scanvic et al., 2023; Chen et al., 2022a; Tachella et al., 2022a). The idea of additional corruption (as we do in our work), is more related to the Noisier2Noise framework (Moran et al., 2020) which has been initially developed for denoising. Millard & Chiew (2023) extended this idea to the MRI setting to learn models for MRI restoration without access to clean data. All these approaches learn a restoration model without access to reference data. However, as they are inherently end-to-end methods, their performance on out-of-distribution tasks (e.g., due to different acquisition trajectories) is known to degrade (Jalal et al., 2021; Zach et al., 2022).
@@ 2.2 Ambient Diffusion for MRI
The MRI acquisition process results in linearly corrupted measurements and thus it should be possible to use the Ambient Diffusion framework to learn from corrupted observations. Yet, we identified three important changes that differ from the setting studied by Daras et al. (2023b): i) the inpainting happens in the Fourier Domain, ii) the inpainting has structure, i.e. whole vertical lines in the spatial Fourier are either observed or not observed (instead of masking random pixels) and, iii) the image is measured from an array of spatially varying receive coils. In what follows, we show how to account for these factors and apply the Ambient Diffusion framework to MRI data.
@@
This is possible to do because we can just subsample the available data ${\bm{z}}_{{\bm{x}},i}$. To further corrupt the given measurements, we keep the inpainting structure that there is in the measurements, i.e. we delete (at random) 2-D lines in Fourier space (instead of pixel masking as in Daras et al. (2023b)). Due to the multiplicity of coils, we combine their measurements to form a crude estimator of ${\bm{x}}$ by taking the adjoint of $\tilde{A}_{\mathrm{train}}$: $\tilde{\bm{y}}_{\mathrm{train}}=\sum_{i}S_{i}^{H}\mathcal{F}^{-1}(\tilde{\bm{z}}_{{\bm{x}},i})$. Notice that for a discrete signal ${\bm{x}}\in\mathbb{C}^{n}$, all these operations are linear and hence $\tilde{\bm{y}}$ can be written as:
@@
Let ${\bm{y}}_{t,\mathrm{train}},\tilde{\bm{y}}_{t,\mathrm{train}}$ represent the noisy versions of ${\bm{y}}_{\mathrm{train}},\tilde{\bm{y}}_{\mathrm{train}}$ respectively, i.e.:
$\begin{cases}{\bm{y}}_{t,\mathrm{train}}=A_{\mathrm{train}}\left({\bm{x}}+\sigma_{t}\bm{\eta}\right)\\ \tilde{\bm{y}}_{t,\mathrm{train}}=\tilde{A}_{\mathrm{train}}\left({\bm{x}}+\sigma_{t}\bm{\eta}\right)\end{cases}.$ (2.9)
@@
The formal proof of this Theorem is given in the Appendix. We provide a sketch of the proof here. The techniques are based on Theorem 4.2 of Ambient Diffusion (Daras et al., 2023b). In Theorem 4.2, the condition that needs to be satisfied is that $\mathbb{E}[A_{\mathrm{train}}^{T}A_{\mathrm{train}}|\tilde{A}_{\mathrm{train}}]$ is full-rank. Our objective is slightly different because the network doesn’t see the whole forward operator $\tilde{A}_{\mathrm{train}}$, but only part of it, i.e. the matrix $\tilde{P}$. By following the steps of the Ambient Diffusion proof, we arrive at the necessary condition for our case, which is to prove that $\mathbb{E}[A_{\mathrm{train}}|\tilde{P}]$ is full-rank.